OpenStudy (anonymous):

Let..... (will draw what to let in a sec!) Which of the following could be a possible decomposition of h(x)? (will post answer choices in a sec also!) ***PLease explain!! Thank youuu :D

OpenStudy (anonymous):



OpenStudy (anonymous):

answer choices A,B,C,D top to bottom :)

jimthompson5910 (jim_thompson5910):

(g o f)(x) means g(f(x))

jimthompson5910 (jim_thompson5910):

basically, you start with g(x) then you replace all copies of 'x' with f(x)

OpenStudy (anonymous):

not quite following :(

jimthompson5910 (jim_thompson5910):

so for example, say f(x) = 3x+1 g(x) = 2x g(f(x)) means g(x) = 2x g(f(x)) = 2*f(x) g(f(x)) = 2(3x+1)

OpenStudy (anonymous):

ohh i see now :)

OpenStudy (anonymous):

so for this equation, we need to solve for f(x) first right?

jimthompson5910 (jim_thompson5910):

for this question, you simply use the steps I did above but for the given functions f(x) and g(x) so for example, in choice D \[\Large g(x) = \frac{x}{x+1}\] \[\Large g(f(x)) = \frac{f(x)}{f(x)+1}\] \[\Large g(f(x)) = \frac{x}{x+1}\] which is NOT what we want, so choice D is out

OpenStudy (anonymous):